[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jan 23, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon Jan 23 16:31:28 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. On Sunday, Jan. 22, Nelsie had problems populating swell1 for the  
Pacific domain from both 06Z data and 12Z data in GFE on the LX3  
workstaiton. No problems occurred on the Atlantic desk workstations. No  
problems were encountered this morning.

2. The latest AWIPS upgrade is now stable. All the high-priority NCF  
tickets have been closed. There are lingering issues with the spell  
checker, banners and the NHA domain (NCF fixed an XML error, will be  
corrected when EDEX server is rebooted later today) that TSB's AWIPS  
SysAdmins Chris Mello and Matt Sardi are following up on.
HSU:
No new issues.
Other items:

NHC RFCs:
1. RFC #826 (OPC) - Send Cryosat-2 and Jason-3 altimeter data to NHC and  
WFO HFO for use in N-AWIPS via DBNet. Implemented on Monday, January 23 at  
1500Z (10 am EST). OPC has shared their table files datatype.tbl and  
miscset.tbl with NHC from their Google drive.

2. RFC #2819 - Make a change to allow the Call Manager to access the NTP  
(Network Time Protocol) server for time synchronization. This is required  
for the Hurricane Hotline to function correctly. To be implemented on  
Monday, January 23 at 1800Z (1 pm EST).

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170123/b80d8bfb/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list